# GW231123: a Binary Black Hole Merger with Total Mass 190-265

### v2

@@ 4.3 Inference
Our Bayesian analysis indicates that GW231123 was produced from a high-mass compact binary merger with highly spinning components. We infer individual source component masses and with spin magnitudes and . We present a summary of the key source properties of GW231123 in Table 2. Unless otherwise stated, we report all mass measurements in the source frame.
@@
[…]
@@
We consistently infer that both BH are highly spinning independent of the model we use. As shown in Figure 4, we infer that the primary spin magnitude at probability and the secondary spin magnitude at probability. The primary component of GW231123 has one of the highest confidently measured BH spins observed through GWs (evidence for highly spinning BH has also been presented in Hannam et al., 2022; Nitz et al., 2020; Abbott et al., 2024a, 2023a; Wadekar et al., 2023; Williams, 2025).
@@
We are unable to reliably infer the spin orientation of the binary; we infer polar angles between each spin vector and the orbital angular momentum that vary not only between models, but also when independently analysing data obtained by LIGO Livingston compared to LIGO Hanford, see Appendix B. To understand these differences, we carried out a series of analyses with different frequency ranges. We found that all models consistently infer greater support for spin components aligned with the orbital angular momentum when independently analysing LIGO Hanford data, and when excluding data from LIGO Livingston below . When spin misalignment is inferred, we are unable to conclusively constrain the spin orientation away from aligned.
@@
The uncertainty in the spin misalignment affects the inferred effective inspiral spin , which parameterizes the spin aligned with the orbital angular momentum (Santamaria et al., 2010; Ajith et al., 2011). Negative would imply that at least one spin is misaligned with the orbital angular momentum by more than ninety degrees. We cannot rule out , but there is an probability that is positive. The inferred effective precessing spin (Schmidt et al., 2015) is consistently measured between models and deviates from the prior, . Although we infer variation between models, we consistently obtain large Bayes factors (5–10) in favour of precession. Since the distribution of Bayes factors from noise alone is unknown, we additionally quantify the evidence for precession in GW231123 by computing the precession SNR, (Fairhurst et al., 2020a, b). In the absence of any precession in the signal, we expect in 90% of cases. We infer an SNR of . Although the high SNR tail is consistent with the large Bayes factors (Green et al., 2021; Pratten et al., 2020b), we infer non-negligible support below . We are therefore unable to confidently claim precession in GW231123. GW190521 was also found to exhibit mild evidence for spin-precession (Abbott et al., 2020d, e).
@@
We observe significant differences in the inferred luminosity distance and inclination angle of GW231123’s source, depending on the model, although most models infer comparable probabilities for face-on vs face-off ( and ) configurations. We also infer substantial variation in the detector-frame quantities, despite seeing agreement between several models in the source-frame parameters. See Appendix A for a detailed discussion. Owing to disagreements in the inferred inclination angle of the binary, we similarly observe differences in the inferred higher-order multipole SNRs from each model. Following the methodology in Abbott et al. (2020f, g), where for each multipole the IMRPhenomXHM signal model (Garc´ıa-Quirós et al., 2020) is used to remove any contribution parallel to the dominant multipole and to calculate the orthogonal optimal SNR (Mills & Fairhurst, 2021), we nevertheless find that all models provide support for the multipole in GW231123. We infer an average orthogonal optimal SNR of when combining the results from all models with equal weight.
@@
The properties of the remnant BH are estimated in different ways depending on the model. We apply the NRSur7dq4Remnant model (Varma et al., 2019) to the samples obtained by NRSur, and we average several fits calibrated to numerical relativity simulations (Hofmann et al., 2016; Healy & Lousto, 2017; Jiménez-Forteza et al., 2017) for samples obtained by v5PHM, TPHM, XPHM, XO4a. When combining the results with equal weight, we infer the final mass and spin of the remnant BH to be and respectively. For certain binary configurations, the remnant BH may receive a recoil velocity that is enough to eject the remnant from its host galaxy (Merritt et al., 2004). We infer a measurement of the remnant BH’s recoil velocity that differs from the effective prior distribution: . This measurement is based on the NRSur analysis and the NRSur7dq4Remnant remnant model, the only fit providing recoil velocities estimates (Varma et al., 2019, 2020).
@@ 4.4 Waveform Consistency Checks
To further assess whether a CBC signal with the inferred parameters in Section 4.3 adequately represents the data, we perform several consistency checks using a signal-agnostic approach that reconstructs coherent transient power, and through a model incorporating a modified wave dispersion relationship. First, we compare the waveform of the maximum-likelihood sample from parameter estimation in Section 4.1 to one obtained through minimally modelled analyses that make no assumptions about the source or morphology of the signal (Szczepańczyk et al., 2021; Salemi et al., 2019; Ghonge et al., 2020). Second, we conduct a residuals analysis, subtracting the best-fit waveform from the detector data and searching for coherent residual power (Abbott et al., 2021b). Discrepancies between the modelled and minimally modelled waveforms, or the presence of significant excess residual power, could indicate physical effects in addition to or alternative to those in our BBH signal models or unaccounted-for noise features (Johnson-McDaniel et al., 2022).
@@
For the waveform reconstruction comparisons, we use BayesWave, cWB-2G, and cWB-BBH for the minimally modelled analysis. To evaluate the agreement between the signal as found by the modelled analysis and minimally modelled approach, we calculate the overlap between the maximum-likelihood sample from parameter estimation using the NRSur model and the median BayesWave or cWB maximum-likelihood waveform. An overlap of 1 indicates perfect agreement, while an overlap of 0 indicates no similarity between waveforms. To assess whether the overlaps are consistent with signals of comparable parameters and noise realizations, we perform a dedicated set of injections wherein we inject waveforms generated by draws from the posterior distribution of the source parameters into detector data surrounding the event. The BayesWave analysis performed 400 injections into approximately 8 hours of data surrounding the event, and the cWB analysis injected about 2800 draws from the posterior distribution in an interval of two weeks around the event. We find good agreement between the minimally modelled and CBC waveform reconstructions. The overlaps between the CBC maximum-likelihood waveform and BayesWave, cWB-2G, and cWB-BBH are 0.97, 0.96, and 0.98, respectively. Compared to the distributions of overlaps from the injections, the -values (defined as the fraction of injections with overlaps below that of the real event) are 0.74, 0.57, and 0.92 for BayesWave, cWB-2G, and cWB-BBH, respectively. Under the null hypothesis, the -values should be distributed uniformly from 0 to 1, so these results indicate that the overlaps are consistent with expectations from systems similar to GW231123.
@@
For the residuals test, we produce residual data by subtracting from the original data the maximum-likelihood waveform from the NRSur parameter estimation samples. If the signal has been modelled sufficiently, the residual data should be consistent with Gaussian noise. We analyze the residual data with BayesWave, and calculate the 90% credible upper limit on the recovered network SNR (). To compare to expected values of , we also analyze segments of data (with no injected signal) selected randomly from 16384 s of data surrounding the event, and calculate the probability of obtaining an higher than that of residual data. Details of this procedure can be found in (Abbott et al., 2021b). We find no significant excess SNR in the residual data beyond what is expected from only Gaussian noise. Compared to the distribution of from the noise-only runs, the -value is 0.35. This further confirms that minimally modelled tests do not flag any features in the data missed by the analyses described in Section 4.1.
@@
We additionally search for post-ringdown echo signals (Tsang et al., 2018, 2020) with a BayesWave-based search, finding negative evidence for their presence (as quantified by the Bayes factor ), consistent with the above findings. As a final consistency check, an analysis incorporating a modified wave dispersion relation due to non-zero graviton mass (Abbott et al., 2021b) yields agreement with massless wave propagation when based on the NRSur or TPHM models. Instead, when assuming the XPHM or XO4a templates, a statistically significant violation is found, suggesting missing signal components not captured by these models, which is consistent with the discussion of systematics in Section 4.2.
@@ 5 Black-hole ringdown
Massive systems dominated by merger-ringdown, such as GW231123, are ideal to test the BH signal interpretation by applying BH spectroscopy techniques (Dreyer et al., 2004; Berti et al., 2009, 2025), yielding remnant properties under minimal assumptions on the remnant formation process. We fit superpositions of damped sinusoids, aiming to associate them with characteristic quasi-normal modes (QNMs) of a BH, which drive its relaxation to equilibrium. The resulting parameter estimates make it possible to robustly validate inspiral–merger–ringdown (IMR) measurements, since a QNM description is generic to any BH remnant (e.g., a BH formed from an eccentric binary).
We truncate portions of data in the time domain at different analysis start times , and fit two sets of models. The first (DS-) is a superposition of damped-sinusoids , with constants as free parameters, assuming (circularly polarized wave). In the second (Kerr), complex frequencies are identified with QNMs of a Kerr BH, and , with detector frame (redshifted) mass , spin , and the QNM angular () and overtone () indices. In addition to the longest-lived , we consider , the linear QNMs with the largest predicted amplitudes for binary mergers (Kamaretsos et al., 2012; London et al., 2014; Cheung et al., 2024; Zhu et al., 2025; Nobili et al., 2025; Carullo, 2024). Here, we include both contributions, accommodating generic signal polarizations, and enters the expression as mode amplitudes are degenerate with the source distance.
We use the pyRing pipeline (Carullo et al., 2019) with standard analysis settings (Isi & Farr, 2021; Abbott et al., 2021b; Gennari et al., 2024). We pre-condition the data by subtracting the Hz power line, which reduces the required analysis duration to s (Siegel et al., 2025). We sample the posterior distribution using the CPNest nested sampling algorithm (Veitch et al., 2020). Times are relative to the median of the polarizations peak time in the LIGO Hanford data, computed from the NRSur reconstructed waveform. The sky location is fixed to the NRSur maximum likelihood value, fixing the analysis start time in LIGO Livingston, as required by the truncated time-domain formulation of the analysis (Isi & Farr, 2021). We have verified that repeating the analysis at different sky location values drawn from the NRSur posterior does not affect our conclusions.
Damped sinusoids are expected to be valid in the stationary regime of BH relaxation, typically (namely ms assuming ) past the signal peak; fitting earlier may provide spurious support for additional modes (Berti et al., 2025). However, a time-domain waveform reconstruction of GW231123 indicates a highly complex morphology, displaying a monotonic decay only after ms, significantly later than the nominal ; see Appendix C. Given this uncertainty, we explore a wide range of times ms in steps of ms.
@@
These multi-mode combinations are in tension with IMR analyses. The most favoured Kerr mode in addition to the according to Bayes factors, the , implies a – distribution that does not overlap with IMR estimates. Adding only results in partial overlap, while the overlaps to a larger degree. However, the short-lived mode alone is not expected to give rise to the features observed in the signal until late times. The mode addition results in the most significant overlap, with a amplitude comparable to the , consistently with later times results. While QNMs can be strongly excited in highly precessing systems with large mass asymmetry (O’Shaughnessy et al., 2013; Zhu et al., 2025; Nobili et al., 2025), the IMR analyses of GW231123 predict minimal power in the mode, as discussed in Appendix D. Highly eccentric configurations can excite modes (Sperhake et al., 2008), but we lack sufficiently complete merger–ringdown models for eccentric-binary signals to reliably assess this possibility; see Section 7 for further discussion. DS- analyses with did not prefer more than two modes, but future investigations will be required to determine whether the observed features can be induced by a superposition of many overlapping modes.
In summary, the Kerr fits recover the remnant spin with large uncertainty, and they robustly predict at all times, supporting the interpretation of a massive BH remnant. Further investigations will be required to characterize the nature of the bi-modal features persistently observed in the signal and consistently interpret the multi-mode fits at earlier times. Given these large uncertainties, we do not consider tests of the no-hair properties of BHs in general relativity that would be enabled by a confident two-mode identification (Dreyer et al., 2004; Berti et al., 2009; Gossan et al., 2012; Brito et al., 2018; Carullo et al., 2018; Bhagwat et al., 2020; Isi & Farr, 2021).

### v3

@@ 4.3 Inference
Our Bayesian analysis indicates that GW231123 was produced from a high-mass compact binary merger with highly spinning components. We infer individual source component masses and with spin magnitudes and . We present a summary of the key source properties of GW231123 in Table 2. Unless otherwise stated, we report all mass measurements in the source frame, and all measurements correspond to the median and 90% symmetric credible level.
@@
Although we observe differences depending on the model, the primary and secondary component masses nevertheless have a significant probability of lying within the mass gap from (pulsational) PISN processes, as shown in Figure 3, where we assume a nominal gap ranging from 60–130 (see Section 6 for a detailed discussion). The binary’s total mass is constrained to be within 189–266 . This measurement exceeds the 95th percentile of the inferred total mass from GW190521 (LIGOScientific:2021vkt). Assuming a FAR threshold of one per year, similar to KAGRA:2021duu, the source of GW231123 is the highest mass BBH observed by the LVK to date; other lower significance high-mass observations have been discussed in LIGOScientific:2021usb; Wadekar:2023gea; Williams:2024tna; Ruiz-Rocha:2025yno.
@@
We consistently infer that both BH are highly spinning independent of the model we use. As shown in Figure 4, we infer that the primary spin magnitude at probability and the secondary spin magnitude at probability, see Sec. 6.5 for details. The primary component of GW231123 has one of the highest confidently measured BH spins observed through GWs (evidence for highly spinning BH has also been presented in Hannam:2021pit; Nitz:2020oeq; LIGOScientific:2021usb; LIGOScientific:2021vkt; Wadekar:2023gea; Williams:2024tna).
@@
We are unable to reliably infer the spin orientation of the binary; we infer polar angles between each spin vector and the orbital angular momentum that vary not only between models, but also when independently analysing data obtained by LIGO Livingston compared to LIGO Hanford, see Appendix B. In an attempt to understand these differences we carried out a series of analyses with different frequency ranges. We found that all models consistently infer greater support for spin components aligned with the orbital angular momentum, and no sign of systematics, when independently analysing LIGO Hanford data, and when excluding data from LIGO Livingston below 50 Hz. However, as with the systematics issues discussed in Section 4.2 and Appendix A, we were not able to conclusively reproduce this behaviour with injections of mock signals, and did not find any significant noise features below that could be the cause, although we did not perform a detailed study of Gaussian noise fluctuations or low-SNR glitches. When spin misalignment is inferred, we are unable to conclusively constrain the spin orientation away from aligned.
@@
The uncertainty in the spin misalignment affects the inferred effective inspiral spin , which parameterizes the spin aligned with the orbital angular momentum (Santamaria:2010yb; Ajith:2009bn). Negative would imply that at least one spin is misaligned with the orbital angular momentum by more than ninety degrees. We cannot rule out , but there is an probability that is positive. The inferred effective precessing spin (Schmidt:2014iyl) is consistently measured between models and deviates from the prior, . Although we infer variation between models, we consistently obtain large Bayes factors () in favor of the precessing hypothesis compared to the spin-aligned hypothesis (spins aligned with the orbital angular momentum). Since the distribution of Bayes factors from noise alone is unknown, we additionally quantify the evidence for precession in GW231123 by computing the precession SNR, (Fairhurst:2019srr; Fairhurst:2019vut). In the absence of any precession in the signal, we expect in 90% of cases. We infer an SNR of . Although the high SNR tail is consistent with the large Bayes factors (Green:2020ptm; Pratten:2020igi), we infer non-negligible support below . We are therefore unable to confidently claim precession in GW231123. GW190521 was also found to exhibit mild evidence for spin-precession (LIGOScientific:2020iuh; LIGOScientific:2020ufj).
@@
We observe significant differences in the inferred luminosity distance and inclination angle of GW231123’s source, depending on the model, although we repeatedly infer nearly symmetric distributions for the inclination angle around for all models except XO4a. We also infer substantial variation in the detector-frame quantities, despite seeing agreement between several models in the source-frame parameters. See Appendix A for a detailed discussion. Owing to disagreements in the inferred inclination angle of the binary, we similarly observe differences in the inferred SNRs in each higher-order multipole obtained by each model. Following the methodology in LIGOScientific:2020stg; LIGOScientific:2020zkf, where for each multipole the IMRPhenomXHM signal model (Garcia-Quiros:2020qpx) is used to remove any contribution parallel to the dominant multipole and to calculate the orthogonal optimal SNR (Mills:2020thr), we nevertheless find that all models provide support for the multipole in GW231123. We infer an average orthogonal optimal SNR of when combining the results from all models with equal weight.
@@
The properties of the remnant BH are estimated in different ways depending on the model. We apply the NRSur7dq4Remnant model (Varma:2019csw) to the samples obtained by NRSur, and we average several fits calibrated to numerical relativity simulations (Hofmann:2016yih; Healy:2016lce; Jimenez-Forteza:2016oae) for samples obtained by v5PHM, TPHM, XPHM, XO4a. When combining the results with equal weight, we infer the final mass and spin of the remnant BH to be and respectively. For certain binary configurations, the remnant BH may receive a recoil velocity that is enough to eject the remnant from its host galaxy (Merritt:2004xa). We infer a measurement of the remnant BH’s recoil velocity that differs from the effective prior distribution: . This measurement is based on the NRSur analysis and the NRSur7dq4Remnant remnant model, the only fit providing recoil velocities estimates (Varma:2019csw; Varma:2020nbm).
@@ 4.4 Waveform Consistency Checks
To further assess whether a CBC signal with the inferred parameters in Section 4.3 adequately represents the data, we perform several consistency checks using a signal-agnostic approach that reconstructs coherent transient power, and through a model incorporating a modified wave dispersion relationship. First, we compare the waveform of the maximum-likelihood sample from parameter estimation in Section 4.1 to one obtained through minimally modelled analyses that make no assumptions about the source or morphology of the signal (Szczepanczyk:2020osv; Salemi:2019uea; Ghonge:2020suv). Second, we conduct a residuals analysis, subtracting the best-fit waveform from the detector data and searching for coherent residual power (LIGOScientific:2021sio). Discrepancies between the modelled and minimally modelled waveforms, or the presence of significant excess residual power, could indicate physical effects in addition to or alternative to those in our BBH signal models or unaccounted-for noise features (Johnson-McDaniel:2021yge).
@@
For the waveform reconstruction comparisons, we use BayesWave (Cornish:2014kda; Cornish:2020dwh; Chatziioannou:2021ezd), cWB-2G (Klimenko:2008fu; Klimenko:2015ypf; Drago:2020kic), and cWB-BBH (Mishra:2024zzs) for the minimally modelled analysis. To evaluate the agreement between the signal as found by the modelled analysis and minimally modelled approach, we calculate the overlap between the maximum-likelihood sample from parameter estimation using the NRSur model and the median BayesWave or cWB maximum-likelihood waveform. An overlap of 1 indicates perfect agreement, while an overlap of 0 indicates no similarity between waveforms. To assess whether the overlaps are consistent with signals of comparable parameters and noise realizations, we perform a dedicated set of injections wherein we inject waveforms generated by draws from the posterior distribution of the source parameters into detector data surrounding the event. The BayesWave analysis performed 400 injections into approximately 8 hours of data surrounding the event, and the cWB analysis injected about 2800 draws from the posterior distribution in an interval of two weeks around the event. We find good agreement between the minimally modelled and CBC waveform reconstructions. The overlaps between the CBC maximum-likelihood waveform and BayesWave, cWB-2G, and cWB-BBH are 0.97, 0.96, and 0.98, respectively. Compared to the distributions of overlaps from the injections, the -values (defined as the fraction of injections with overlaps below that of the real event) are 0.74, 0.57, and 0.92 for BayesWave, cWB-2G, and cWB-BBH, respectively. Under the hypothesis that the overlaps between the BayesWave and cWB reconstructions with the maximum-likelihood waveform are drawn from the same distribution as the injection overlaps, the -values should be distributed uniformly from 0 to 1, so these results indicate that the overlaps are consistent with expectations from systems similar to GW231123.
@@
For the residuals test, we produce residual data by subtracting from the original data the maximum-likelihood waveform from the NRSur parameter estimation samples. If the signal has been modelled sufficiently, the residual data should be consistent with Gaussian noise. We analyze the residual data with BayesWave, and calculate the 90% credible upper limit on the recovered network SNR (). To compare to expected values of , we also analyze segments of data (with no injected signal) selected randomly from 16384 s of data surrounding the event, and calculate the probability of obtaining an higher than that of residual data. Details of this procedure can be found in (LIGOScientific:2021sio). We find no significant excess SNR in the residual data beyond what is expected from only Gaussian noise. Compared to the distribution of from the noise-only runs, the -value is 0.35. This further confirms that minimally modelled tests do not flag any features in the data missed by the analyses described in Section 4.1.
@@
We additionally search for post-ringdown echo signals (Tsang:2018uie; Tsang:2019zra) with a BayesWave-based search, finding negative evidence for their presence (as quantified by the Bayes factor ), consistent with the above findings. As a final consistency check, an analysis incorporating a modified wave dispersion relation due to non-zero graviton mass (LIGOScientific:2021sio) yields agreement with massless wave propagation when based on the NRSur or TPHM models. Instead, when assuming the XPHM or XO4a templates, a statistically significant violation is found, suggesting missing signal components not captured by these models, which is consistent with the discussion of systematics in Section 4.2.
@@ 5 Black-hole ringdown
Massive systems dominated by merger-ringdown, such as GW231123, are ideal to test the BH signal interpretation by applying BH spectroscopy techniques (Detweiler:1980gk; Dreyer:2003bv; Berti:2005ys; Berti:2009kk; Berti:2025hly), yielding remnant properties under minimal assumptions on the remnant formation process. We fit superpositions of damped sinusoids, aiming to associate them with characteristic quasi-normal modes (QNMs) of a BH, which drive its relaxation to equilibrium. In principle, the resulting parameter estimates make it possible to robustly validate IMR measurements, since a QNM description is generic to any BH remnant (e.g., a BH formed from an eccentric binary).
We truncate portions of data in the time domain at different analysis start times , and fit two sets of models. The first (DS-) is a superposition of damped-sinusoids , with constants as free parameters, assuming (circularly polarized wave). In the second (Kerr), complex frequencies are identified with QNMs of a Kerr BH, and , with detector frame (redshifted) mass , spin , and the QNM angular () and overtone () indices. In addition to the longest-lived , we consider , the linear QNMs with the largest predicted amplitudes for binary mergers (Kamaretsos:2012bs; London:2014cma; Cheung:2023vki; Zhu:2023fnf; Nobili:2025ydt; Carullo:2024smg). Here, we include both contributions, accommodating generic signal polarizations, and enters the expression as mode amplitudes are degenerate with the source distance.
We use the pyRing pipeline (Carullo:2019flw) with standard analysis settings (Isi:2021iql; LIGOScientific:2021sio; Gennari:2023gmx). We pre-condition the data by subtracting the Hz power line, which reduces the required analysis duration to s (Siegel:2024jqd). We sample the posterior distribution using the CPNest nested sampling algorithm (cpnest099). Times are relative to s in the LIGO Hanford data, compatible with the median of the polarizations peak time from the NRSur reconstructed waveform, subject to an uncertainty s. The sky location is fixed to a value compatible with the NRSur maximum likelihood value, fixing the analysis start time in LIGO Livingston, as required by the truncated time-domain formulation of the analysis (Isi:2021iql). We have verified that repeating the analysis at different sky location values drawn from the NRSur posterior does not affect our conclusions.
Damped sinusoids are expected to be valid in the stationary regime of BH relaxation, typically (namely ms assuming ) past the signal peak; fitting earlier may provide spurious support for additional modes (Berti:2025hly). However, a time-domain waveform reconstruction of GW231123 indicates a highly complex morphology, displaying a monotonic decay only after ms. This is significantly later than the nominal , after which monotonic decay is expected for vanilla signal morphologies; see Appendix C. Given this uncertainty, we explore a wide range of times ms in steps of ms.
@@
These multi-mode combinations are in tension with IMR analyses. The most favoured Kerr mode in addition to the according to Bayes factors, the , implies a – distribution that does not overlap with IMR estimates. Adding only results in partial overlap, while the overlaps to a larger degree. However, the short-lived mode alone is not expected to give rise to the features observed in the signal until late times. The mode addition results in the most significant overlap, with a amplitude comparable to the , consistently with later times results. While QNMs can be strongly excited in highly precessing systems with large mass asymmetry (OShaughnessy:2012iol; Zhu:2023fnf; Nobili:2025ydt), the IMR analyses of GW231123 predict minimal power in the mode, as discussed in Appendix D. Highly eccentric configurations can excite modes (Sperhake:2007gu), but we lack sufficiently complete merger–ringdown models for eccentric-binary signals to reliably assess this possibility; see Section 7 for further discussion. DS- analyses with did not prefer more than two modes, but future investigations will be required to determine whether the observed features can be induced by a superposition of many overlapping modes.
In summary, the Kerr fits recover the remnant spin with large uncertainty, and they robustly predict at all times, supporting the interpretation of a massive BH remnant. Further investigations will be required to characterize the nature of the bi-modal features persistently observed in the signal and consistently interpret the multi-mode fits at earlier times. Given these large uncertainties, we do not consider tests of the no-hair properties of BHs in general relativity that would be enabled by a confident two-mode identification (Detweiler:1980gk; Dreyer:2003bv; Berti:2005ys; Berti:2009kk; Gossan:2011ha; Brito:2018rfr; Carullo:2018sfu; Bhagwat:2019bwv; Isi:2021iql; Berti:2025hly).